OpenStudy (luigi0210):

Find the domain of \(f(x)=sin^{-1}(x-1)\)

OpenStudy (anonymous):

the domain of \(\sin^{-1}(x)\) is \([-1,1]\) so your job is to solve \(x-1=-1\) for the left hand endpoint and \(x-1=1\) for the right hand endpoint